OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6? Giving metal to the right answer.

OpenStudy (anonymous):

help please

OpenStudy (shamil98):

parallel lines have equal slopes, so first change your equation to slope-intercept form. y = mx + b then input your given coordinates and the slope (m) into the point-slope formula to get an equation. y - y1 = m (x-x1)